OpenStudy (vera_ewing):

The following galvanic cell at standard conditions has a potential of 0.03 V: Ag+(aq) + Fe2+(aq) → Fe3+(aq) + Ag(s). What would be the effect of increasing the concentration of Ag+? A. The cell potential will be 0 V. B. The cell potential will decrease. C. The cell potential will not change. D. The cell potential will increase.

OpenStudy (vera_ewing):

@nechirwan Help!

OpenStudy (aaronq):

Use the Nernst equation: \(\xi=\xi^o+\dfrac{RT}{nF}ln(\dfrac{[Fe^{2+}]}{[Ag^+]}{}) \) At 25 celsius: \(\xi=\xi^o+\dfrac{0.0592}{n}log(\dfrac{[Fe^{2+}]}{[Ag^+]}{}) \) you basically wanna compare the values you get by changing the concentration, so At std cond, \([Fe^{2+}]=[Ag^+]=1\), therefore \(log\dfrac{1}{1}=0\) the cell potential is 0.03V When \([Ag^+]\) is increased, \(log\dfrac{1}{2}=-0.3010\) if you work it into the equation you'll notice the voltage increased.

OpenStudy (aaronq):

by the way i messed up the second equation, it should be: \(\xi =\xi^o-\dfrac{0.0592}{n}log\dfrac{[Fe^{2+}]}{[Ag^+]}\)